What concentration of HCl will have a pH of 5.0?

b) What concentration of acetic acid will have a pH of 5.0? (For acetic acid, pKa=4.76)

c) What concentration of phosphoric acid will have a pH of 5.0? (For H3PO4, pKa1 = 2.12, pKa2 = 7.21, pKa3 =12.32)

d) What is the pH of a 10 mM solution of phosphoric acid?

e) How much 1.0 M NaOH must be added to 100 ml of 10 mM H3PO4 to raise the pH of the solution to 7.0?

f) If 2 volumes of 0.2 M Na2HPO4 are added to 1 volume of 0.1 M H3PO4, what is the pH of the resulting phosphate buffer?
Answer #1

(a) HCl is strong acid . [H+] = [HCl]

pH = 5.0 = -log[H+]

[HCl] = 10-5.0 = 1.0×10-5M (answer)

(b) acetic acid is a weak acid, for weak acid .

Ka = 10-4.76 = 1.74×10-5 = [H+] × \alpha/(1-\alpha) = 1.0×10-5× \alpha/(1-\alpha)

\alpha = 1.74/2.74 = 0.63474

[H+] = c × \alpha

c = 1.0×10-5 /0.63474 = 1.575×10-5 M

Concentration of acetic acid = 1.58×10-5 M. (Answer)

(c) phosphoric acid is weak acid and 2nd dissociation constant is very small. We can consider concentration of H+ from 1st dissociation of acid only.

Ka1= 10-2.12 = 1.0×10-5(\alpha /1-\alpha)

\alpha = 0.9987

[H+] = c \alpha

Concentration of phosphoric acid = 1.0×10-5/0.9987

= 1.0×10-5 M. (Answer)

(d)

c = 10mM = 1×10-2M

Ka1 = c\alpha2/(1-\alpha)

\alpha2/(1-\alpha) = 0.7586

\alpha = 0.57068

[H+] = c \alpha = 1×10-2 × 0.57068 M

pH = -log(5.7068×10-3) = 2.244

pH = 2.24. (Answer)